OpenStudy (anonymous):

Determine function g such that h(x) = f(g(x)). h(x) = x^8 + 8 f(x) = x + 8 g(x)= ?

OpenStudy (swissgirl):

alrighty so we know that our f(x)=x+8

OpenStudy (swissgirl):

ill let akitav take over

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

It is given that for x, f(x) = x+8, so f(g(x))= g(x) +8. Since,h(x)=f(g(x)) Put the value of h(x) and f(g(x)) here\[x^8+8 = g(x) +8\]Now solve for g(x) as you would any ordinary equation for one variable, and you'd have the expression for g(x).

OpenStudy (anonymous):

so ist it x^8+0 or x^8?

OpenStudy (anonymous):

correct\[g(x)=x^8\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You are welcome :)